# Simpler Hamiltonian structure of Gravity in degenerate limit

### v3

@@ III A different degenerate limit: Hamiltonian structure without ordering ambiguity
The triad transforms as earlier, whereas the gauge field now transform differently. Note that this does not preserve the symplectic form. Let us also demand that the $E^{a}_{i}\partial_{a}$ be preserved (instead of the tetrad one-form), implying that: $\partial_{a}=\frac{\delta_{(a)}}{\delta}\partial^{\prime}_{a}$.
@@
$+$ $E^{{}^{\prime}b}_{i}A_{b}^{{}^{\prime}i}\frac{\delta}{\delta_{(a)}}\frac{\delta}{\delta_{(b)}}\partial_{a}\delta_{(b)}-E^{{}^{\prime}b}_{i}A_{a}^{{}^{\prime}i}\frac{\delta}{\delta_{(b)}}\frac{\delta}{\delta_{(a)}}\partial_{b}\delta_{(a)}\Big{]},$
$H$ $=$ $\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}A_{a}^{{}^{\prime}i}A_{b}^{{}^{\prime}j}~{}+~{}(1+\eta^{2})\delta^{2}\Big{(}\frac{1}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}+\omega_{a}^{{}^{\prime}ik}\omega_{b}^{{}^{\prime}kj}\Big{)}~{}+~{}(1+\eta^{2})\omega^{{}^{\prime}ij}_{b}\frac{\delta^{2}}{\delta_{(a)}\delta_{(b)}}\partial_{a}\frac{\delta_{(b)}}{\delta}\Big{]}$
@@
$H^{\prime}$ $=$ $-\frac{\kappa(1+\eta^{2})}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\bar{R}_{ab}^{{}^{\prime}ij}(E^{\prime}),$
where $D^{\prime}_{a}E^{{}^{\prime}a}_{i}=\partial_{a}E^{a}_{i}$ and $\bar{R}_{ab}^{{}^{\prime}ij}(E^{\prime})=\partial_{[a}\omega_{b]}^{{}^{\prime}ij}(E^{\prime})+\omega_{[a}^{{}^{\prime}ik}(E^{\prime})\omega_{b]}^{{}^{\prime}kj}(E^{\prime})$ is the spatial three-curvature. In obtaining the above, we have redefined the Lagrange multipliers as: $\omega_{t}^{~{}ij}=\frac{1}{\delta}\omega_{t}^{{}^{\prime}ij},~{}N^{a}=\frac{\delta}{\delta_{(a)}}N^{{}^{\prime}a}$.
This reflects another intriguing limit, where the diffeomorphism becomes trivial as in the earlier section. The rotation and Hamiltonian constraints above depend only on the conjugate momenta $E^{a}_{i}$. The quantization of this theory would be free of any operator ordering ambiguity problem.
IV A Connection between degenerate and Carroll limit
In this section, we demonstrate that the degenerate limit of Hilbert-Palatini gravity contains Carroll gravity as a special case. The latter had originally been formulated as a truncated Hamiltonian theory in the limit $c\rightarrow 0$. Recently, both the ‘electric’ and ‘magnetic’ forms of this limit was shown to originate from the Hilbert-Palatini action, and the associated connection-based Hamiltonian formulations were also presented [15].
Using the same parametrization as earlier, the Hilbert-Palatini constraints are given by:
@@ IV A Connection between degenerate and Carroll limit
$H~{}$ $=$ $~{}-~{}\frac{\kappa}{2\sqrt{E}}E^{[a}_{i}E^{b]}_{j}\Big{[}K_{a}^{i}K_{b}^{j}~{}+~{}\partial_{a}\omega_{b}^{~{}ij}(E)+\omega_{b}^{~{}ik}(E)\omega_{b}^{~{}kj}(E)\Big{]}$ (13)
where $K_{a}^{i}:=\omega_{a}^{~{}0i}$. The analogue of the scaling transformation (7) on conjugate pairs is given by:
@@
Note that the transformation of the derivatives and of $\omega_{a}^{{}^{\prime}ij}(E)$ remain the same as in section-II. Let us now find the reduced set of constraints in the degenerate limit, where the scaling transformation is global: $\delta_{(a)}\equiv const.$
IV.1 Electric Carroll gravity
Under (14), the rotation constraint under the scaling remain form-invariant:
@@ IV.1 Electric Carroll gravity
Notice that the Hamiltonian structure above corresponds precisely to the ‘electric’ Carroll limit in connection formulation [15]. The additional factor $\frac{1}{\sqrt{E^{\prime}}}$ in the Hamiltonian constraint $H$ here compared to ref.[15] originates due to the use of a scalar lapse.
@@ IV.2 Magnetic Carroll gravity
Next, we consider a global scaling analogous to eq.(12), which provides another way to implement the degenerate limit:
@@
Using the same redefinition of the Lagrange multipliers as in sec-III, we find that the rotation and spatial diffeomorphism constraints remain form-invariant, whereas the Hamiltonian constraint reduces to:
$H^{\prime}$ $=$ $-\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}\partial^{\prime}_{a}\omega_{b}^{{}^{\prime}ij}(E^{\prime})+\omega_{a}^{{}^{\prime}ik}(E^{\prime})\omega_{b}^{{}^{\prime}kj}(E^{\prime})\Big{]}$
@@ V Conclusions
Here we have constructed a limit of Hamiltonian gravity theory corresponding to vanishing densitized triad. This is equivalent to the limit to a vanishing metric determinant. This limit is relevant in the context of connection-triad formulation of canonical quantum gravity.
Here we unravel two different limits of the Barbero-Immirzi formulation. In both cases, the spatial diffeomorphism constraints are satisfied trivially. Further, the respective Hamiltonian constraints exhibit significant simplifications. In the case of the first limit, an ultralocal behaviour (absence of space derivatives) emerges. A similar behaviour is also encountered in other contexts, for instance, in studies of the BKL limit to singularity. However, let us also emphasize that the associated Hamiltonian structure here is quite different from the one corresponding to a BKL behaviour [23, *sloan1]. A canonical quantization of this degenerate limit may be taken up based on either the triad or connection representation. For the second limit, a triad representation would be more appropriate since the constraints in this case depend only on the densitized triad and its derivatives. It could be interesting to explore the possible (kinematical) consequences for quantum geometry, e.g. the area operator in such limits. A study of the possible perturbative theories around this limit could also be worthwhile.
Our analysis also provides a different interpretation to Carroll gravity in terms of the spatial metric determinant. This may be contrasted with the original interpretation in terms of the $c\rightarrow 0$ limit, or a strong gravitational coupling. The question as to which one is to be preferred is purely optional.
@@ Acknowledgements.
Appendix A A Corollary on the violation of Birkhoff’s theorem in electric Carroll gravity

### v4

@@ III An alternative limit: Hamiltonian structure without ordering ambiguity
To compare with eq.(7), the triad fields transform as earlier, whereas the gauge fields now transform differently. Note that this does not preserve the symplectic form. Let us also demand that $(E^{a}_{i}\partial_{a})$ be preserved (instead of the tetrad one-form), implying that: $\partial_{a}=\frac{\delta_{(a)}}{\delta}\partial^{\prime}_{a}$.
@@
$+$ $E^{{}^{\prime}b}_{i}A_{b}^{{}^{\prime}i}\frac{\delta}{\delta_{(a)}}\frac{\delta}{\delta_{(b)}}\partial_{a}\delta_{(b)}-E^{{}^{\prime}b}_{i}A_{a}^{{}^{\prime}i}\frac{\delta}{\delta_{(b)}}\frac{\delta}{\delta_{(a)}}\partial_{b}\delta_{(a)}+\Big{(}\eta+\frac{1}{\eta}\Big{)}\delta\Bigg{(}A_{a}^{{}^{\prime}i}+\frac{\eta}{2}\epsilon^{ijk}\omega_{a}^{{}^{\prime}jk}\Bigg{)}G^{{}^{\prime}rot}_{i}\Big{]},$
$H$ $=$ $\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}A_{a}^{{}^{\prime}i}A_{b}^{{}^{\prime}j}~+~(1+\eta^{2})\delta^{2}\Big{(}\frac{1}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}+\omega_{a}^{{}^{\prime}ik}\omega_{b}^{{}^{\prime}kj}\Big{)}~+~(1+\eta^{2})\omega^{{}^{\prime}ij}_{b}\frac{\delta^{2}}{\delta_{(a)}\delta_{(b)}}\partial_{a}\frac{\delta_{(b)}}{\delta}\Big{]}$ (15)
@@
$H^{\prime}$ $=$ $-\frac{\kappa(1+\eta^{2})}{2\sqrt{E^{\prime}}}E^{{}^{\prime}a}_{i}E^{{}^{\prime}b}_{j}\bar{R}_{ab}^{{}^{\prime}ij}(E^{\prime}),$ (16)
where $D^{\prime}_{a}E^{{}^{\prime}a}_{i}=\partial^{\prime}_{a}E^{{}^{\prime}a}_{i}$ and $\bar{R}_{ab}^{{}^{\prime}ij}(E^{\prime})=\partial_{[a}\omega_{b]}^{{}^{\prime}ij}(E^{\prime})+\omega_{[a}^{{}^{\prime}ik}(E^{\prime})\omega_{b]}^{{}^{\prime}kj}(E^{\prime})$ is the spatial three-curvature. In obtaining the above, we have redefined the Lagrange multipliers as: $\omega_{t}^{~ij}=\frac{1}{\delta}\omega_{t}^{{}^{\prime}ij},~N^{a}=\frac{\delta}{\delta_{(a)}}N^{{}^{\prime}a}$.
This reflects another intriguing limit, where the diffeomorphism becomes trivial as in the earlier section. However, the rotation and Hamiltonian constraints above depend only on the conjugate momenta $E^{a}_{i}$. Since these have no variables with nontrivial Poisson brackets among themselves, the quantum constraints would have no non-commuting operators. In other words, the quantization of this theory would be free of any operator ordering ambiguity problem, even though regularization issues associated with product of commuting operators would still be alive.
IV Degenerate limit and Carroll gravity
In this section, we demonstrate that the degenerate limit of Hilbert-Palatini gravity reproduce Carroll gravity when the scaling transformations parametrizing the vanishing of the determinant are made global: $\delta_{(a)}\equiv const$. Carroll gravity had originally been formulated as a truncated Hamiltonian theory in the limit $c\rightarrow 0$ [29]. Recently, both the ‘electric’ and ‘magnetic’ forms of this formulation was shown to originate from the Hilbert-Palatini action corresponding to first-order gravity, and the associated connection-based Hamiltonian formulations were also presented [30].
Using the same parametrization of the variables as earlier, the Hilbert-Palatini constraints are given by:
@@ IV Degenerate limit and Carroll gravity
$H$~ $=$ $~-~\frac{\kappa}{2\sqrt{E}}E^{[a}_{i}E^{b]}_{j}\Big{[}K_{a}^{i}K_{b}^{j}~+~\partial_{a}\omega_{b}^{~ij}(E)+\omega_{a}^{~ik}(E)\omega_{b}^{~kj}(E)\Big{]}$ (17)
where $K_{a}^{i}:=\omega_{a}^{~0i}$. We now proceed to find the reduced set of constraints in the degenerate limit implemented through a global scaling transformation.
IV.1 Electric Carroll gravity
First we consider the analogue of the scaling transformation (7) on conjugate pairs is given by:
@@ IV.1 Electric Carroll gravity
Note that the transformation of the derivatives and of $\omega_{a}^{{}^{\prime}ij}(E)$ remain the same as in section-II. Under (18), the rotation constraint under the scaling remain form-invariant:
@@
Notice that the Hamiltonian structure above corresponds precisely to the ‘electric’ Carroll limit in connection formulation [30]. The additional factor $\frac{1}{\sqrt{E^{\prime}}}$ in the Hamiltonian constraint $H$ here compared to ref.[30] originates due to the use of a scalar lapse as opposed to a densitized one.
@@ IV.2 Magnetic Carroll gravity
Next, we consider a global scaling analogous to eq.(14), which provides an alternative realization of the degenerate limit:
@@
Using the redefinition of the Lagrange multipliers as: $\omega_{t}^{~ij}=\frac{1}{\delta}\omega_{t}^{{}^{\prime}ij},~N^{a}=\frac{1}{\delta_{(a)}}N^{{}^{\prime}a},~N=\delta N^{\prime}$, we find that the rotation and spatial diffeomorphism constraints remain form-invariant, whereas the Hamiltonian constraint reduces to:
$H^{\prime}$ $=$ $-\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}\partial^{\prime}_{a}\omega_{b}^{{}^{\prime}ij}(E^{\prime})+\omega_{a}^{{}^{\prime}ik}(E^{\prime})\omega_{b}^{{}^{\prime}kj}(E^{\prime})\Big{]}$ (23)
@@
Note that the ‘electric’ and ‘magnetic’ Carroll limits could be obtained through a $c\rightarrow 0$ limit, which could in turn be defined through a set of global transformations on the tetrad and connection variables leading to a singular determinant, as explicitly demonstrated in Ref.[30]. While the transformations here are global as well, they act on the canonical variables in a different way.
@@ V Conclusions
Here we have presented a Hamiltonian form of gravity theory in the limit of a vanishing triad determinant. This is equivalent to the limit to a vanishing metric determinant. This limit is relevant in the context of connection-triad formulation of canonical quantum gravity. Cases of spacelike singularities (e.g. the cosmological singularity) provide a more specific context where such a limit typically shows up.
We have explored the implications of the possible degenerate limits within the Barbero-Immirzi formulation, which provides a modern perspective to canonical gravity. We elucidate two possible realizations, and find that in both cases the constraint structure is significantly simplified compared to standard Hilbert-Palatini gravity. In particular, the spatial diffeomorphism constraints are shown to be solved trivially, while the Hamiltonian constraint exhibits intriguing simplifications. In the first case, the Hamiltonian constraint becomes a polynomial function of the canonical variables (upto an overall density factor), and in the second it becomes free of ordering ambiguity while being non-polynomial. Either cases provide simpler starting points for a canonical quantization programme when compared to the Hilbert-Palatini Hamiltonian theory. To emphasize, our analysis leads to two main insights regarding the open questions involving the diffeomorphism and Hamiltonian constraints within a canonical quantization approach. Firstly, we learn that the spatial diffeomorphism constraints could be satisfied classically in any physical scenario (e.g. as a spacelike singularity is approached) where the spatial metric tends to degenerate. Hence, it might not be necessary in this regime to carry the diffeomorphisms all the way to the quantum theory, invoke its operator representation and then find its solutions unlike the usual approach. Secondly, the simplification of the Hamiltonian constraint is not an exclusive feature of Ashtekar’s self-dual gravity with complex variables or of Carrollian gravity typically interpreted as the low speed of light limit.
We have also shown that in the first of the two degenerate limits, an ultralocal behaviour (absence of space derivatives) emerges. While a similar behaviour is encountered in some other contexts, for instance, in the context of the BKL limit to a cosmological singularity, let us emphasize that the associated Hamiltonian structure here is quite different from the BKL Hamiltonian formulation [37, *sloan1]. This earlier work is built upon a set of densitized spatial derivatives, whose action is defined in such a way that leads to a set of truncated constraints. Unlike our case, this set contains the spatial diffeomorphism constraints. We also note that ultralocality has been reported in the earlier literature in contexts such as electric Carrollian gravity [8] or self-dual Euclidean gravity in the strong coupling limit [11].
Classically, the fact that the set of constraints in the degenerate limit is weaker suggests that the associated ultralocal behaviour is potentially associated with a higher number of degrees of freedom. The true nature and role of these additional degrees of freedom need a deeper study.
Regarding the possible method to proceed to a canonical quantization, the degenerate limit in Sec-II could be taken up based on either the triad or connection representation. For the alternative limit discussed in Sec-III, a triad representation would be more appropriate since all the constraints in this case depend only on the densitized triad and its derivatives. It might also be worthwhile to explore if a perturbative theory of gravity could be defined around the degenerate limit.
Finally, our analysis provides a different interpretation to Carroll gravity in terms of the spatial metric determinant. Note that the original Carrollian formulations defined in terms of dimensionful constants, i.e. the $c\rightarrow 0$ or $G\rightarrow\infty$ limit, are severely dependent on the choice of units. The framework here is manifestly free of such ambiguities. Further, given the connection between Carrollian physics and null hypersurfaces (e.g. black hole event horizons) [39], it would be interesting to explore a potential role of a degenerate metric limit in such contexts. As our results in Appendix B reflect, the spherically symmetric sector admits solutions which evade the Birkhoff’s theorem in the degenerate limit (as well as in electric Carroll gravity). It could be worthwhile to understand the implications in the broad context of black hole singularities.
@@ Acknowledgements.
Appendix A Derivation of the symplectic form
We begin with the standard parametrization of the tetrad fields in terms of the lapse function and shift vector [15, 35, 34]:
$e^{I}_{t}=NM^{I}+N^{a}V_{a}^{I},~e^{I}_{a}=V^{I}_{a};$
$M_{I}V_{a}^{I}=0,~M_{I}M^{I}=-1;$
$e^{t}_{I}=-\frac{M_{I}}{N},~e^{a}_{I}~=~V^{a}_{I}+\frac{N^{a}M_{I}}{N}~~;$
$M^{I}V_{I}^{a}:=0,~V_{a}^{I}V^{b}_{I}~:=~\delta_{a}^{b},~V_{a}^{I}V^{a}_{J}:=\delta^{I}_{J}+M^{I}M_{J}~~;$
$q:=\det q_{ab}=\det(V_{a}^{I}V_{bI})$
Using this decomposition, the Lagrangian density could be rewritten as:
${\cal L}(e,\omega)$ $=$ $\frac{1}{2\kappa}\Big{[}ee^{\mu}_{I}e^{\nu}_{J}R_{\mu\nu}^{~~~IJ}(\omega)~+~\eta\epsilon^{\mu\nu\alpha\beta}\Big{(}D_{\mu}e_{\nu}^{I}D_{\alpha}e_{\beta I}-e_{\mu}^{I}e_{\nu}^{J}R_{\alpha\beta IJ}(\omega)\Big{)}\Big{]}$ (24)
$=$ $\frac{1}{2}\pi^{a}_{~IJ}\partial_{t}\omega_{a}^{(\eta)IJ}-\frac{1}{2}\omega_{t}^{~IJ}G_{IJ}-N^{a}H_{a}-NH$
where the momenta conjugate to the canonical coordinates $\omega_{a}^{(\eta)IJ}$ are defined as: $\pi^{a}_{~IJ}=ee^{t}_{[I}e^{a}_{J]}$. Since the variables $\omega_{t}^{~IJ},~N^{a}$ and $N$ have no conjugate momenta, these lead to secondary constraints given by the following expressions, respectively:
$G_{IJ}$ $=$ $-D_{a}\pi^{(\eta)a}_{IJ}a\approx 0,$
$H_{a}$ $=$ $\frac{1}{2}\pi^{b}_{~IJ}R^{(\eta)IJ}_{ab}\approx 0,$
$H$ $=$ $\frac{1}{2\sqrt{q}}\pi^{a}_{~IK}\pi^{b}_{~JL}\eta^{KL}R^{(\eta)IJ}_{ab}\approx 0,$ (25)
where $\eta^{IJ}$ is the internal Minkowski metric. From eq.(24), the symplectic form could be read off as: $\Omega=\frac{1}{2}\pi^{a}_{~IJ}\partial_{t}\omega_{a}^{(\eta)IJ}$.
The original canonical momenta $\pi^{a}_{~IJ}$ are not all independent, as these are associated with the (set of six) primary constraints:
$\epsilon^{IJKL}\pi^{a}_{~IJ}\pi^{b}_{~KL}\approx 0$ (26)
These constraints are solved [35] by the following choice of variables, eliminating six redundant components out of the nine components of $\pi^{a}_{~ij}$, leaving only three independent ones ($\chi_{i}$):
$\pi^{a}_{~ij}=\chi_{[i}\pi^{a}_{~0j]}$ (27)
The symplectric form becomes:
$\Omega=\frac{1}{2}\pi^{a}_{~IJ}\partial_{t}\omega_{a}^{(\eta)IJ}=\pi^{a}_{~0i}\partial_{t}\omega_{a}^{(\eta)0i}+\frac{1}{2}\pi^{a}_{~ij}\partial_{t}\omega_{a}^{(\eta)ij}=E^{a}_{i}\partial_{t}A_{a}^{i}+\zeta^{i}\partial_{t}\chi_{i}$ (28)
where we have introduced the new variables as: $E^{a}_{i}:=\pi^{a}_{~0i},~A_{a}^{i}:=\omega_{a}^{(\eta)0i}-\chi_{j}\omega_{a}^{(\eta)ij},\zeta^{i}=-\pi^{a}_{j}\omega_{a}^{(\eta)ij}$. Note that $q=\det E^{a}_{i}:=E$. In the above, the L.H.S. contains eighteen canonical pairs, whereas the R.H.S. has only twelve. Thus, the six redundant components of $\omega_{a}^{~ij}$ have disappeared.
Next, in the time-gauge, $\chi_{i}=0$. With this, the symplectic form (28) finally reduces to:
$\Omega=E^{a}_{i}\partial_{t}A_{a}^{i}$ (29)
This leaves $(A_{a}^{i},E^{b}_{j})$ as the only independent canonical pair subject to the constraints.
Appendix B A Corollary on the violation of Birkhoff’s theorem in electric Carroll gravity